OpenStudy (anonymous):

how do I solve x/-2-3=-5

OpenStudy (anonymous):

x/(-2-3)=-5 x=-5(-2-3) x=10+15 x=25

OpenStudy (anonymous):

or (x/-2)-3=-5 x/-2=-2 x=4 Parenthesis are important in math!